Given that T = 28.3 kN and W = 25 kN, determine the magnitude and sense of t he moments about point B of the following: a. the force T; b. the force W; c. forces T and W combined.
1
Expert's answer
2021-09-07T16:31:26-0400


a) Moment due to force T about B "= T cos 45° \times 20"

"= 28.3 cos 45° \times 20 \\

= 297.33 \;kNm"

Positive and anticlockwise

b) Moment due to the W about B "= -W \times 16"

"= -25 \times 16 \\

= -400 \;kNm"

Negative and clockwise

c) Moment due to T and W about B "= 297.33-400"

"= -102.67 \; kNm"

Negative and clockwise